An evaporator is used to concentrate 12 000 kg/h of a 15% solution of NaOH entering at 60oC to a product of 50% solid. The pressure of the saturated steam used is 300 kPa and the pressure in the vapour space of the evaporator is 11.7 kPa. The overall heat transfer coefficient is 1 560 W/m2K. Calculate the steam used and the steam economy in kg vaporised/kg steam.


1
Expert's answer
2021-10-10T09:32:10-0400

Given:

F = 4536 kg/h

TF = 60 ºC

xF = 0.2

S = ?

Solution


  For the total balance,

        "F = 4536 = L + V"

  For the balance on the solute alone,

    "F \times F = L \times L"

        "4536 (0.2) = L (0.5)"

            L = 1814 kg/h of liquid

  Substituting into total balance and solving,

            V = 2722 kg/h of vapor


To determine T1 = Tsat + BPR of the 50 % concentrate product, first we obtain Tsat of pure water from steam table. At 11.7 kPa,

Tsat = 48.9 ºC.